OpenStudy (anonymous):

the random variable x has a probability distribution of x=2,3,5,7,11 and p(x)=1/6,1/3,1/4,x and y Find the numerical value of x and y and

OpenStudy (anonymous):

if \[\sum(x)=14/3\]

OpenStudy (anonymous):

Are you asking if the sum of the x's is=14/3 and the values of x=2,3,5,7,11, what are the values of x,y? when the P(x) {output value}for each x is 1/6,1/3,1/4,x,y respectively?

OpenStudy (anonymous):

yap